OpenStudy (anonymous):

Find dy/dx for y=ln(5-x)^6

OpenStudy (campbell_st):

rewrite the question using log laws and it becomes \[y = 6\ln(x - 5)\] now you can use the product rule to find the derivative

OpenStudy (anonymous):

y=ln(5-x)^6 y'=6ln(5-x)^5 × -1 = -6ln(5-x)^5